# A phenomenological gravitational waveform model of binary black holes incorporating horizon fluxes

### v2

@@ Abstract
Subjected to the tidal field of its companion, each component of a coalescing binary suffers a slow change in its mass (tidal heating) and spin (tidal torquing) during the inspiral and merger. For black holes, these changes are associated with their absorption of energy and angular momentum fluxes. This effect modifies the inspiral rate of the binary, and consequently, the phase and amplitude of its gravitational waveform. Numerical relativity waveforms contain these effects inherently, whereas analytical approximants for the early inspiral phase have to include them manually in the energy balance equation. In this work, we construct a frequency-domain gravitational waveform model that incorporates the effects of tidal heating of black holes. This is achieved by recalibrating the inspiral phase of the waveform model IMRPhenomD to incorporate the phase corrections for tidal heating. We also include corrections to the amplitude, but add them directly to the inspiral amplitude model of IMRPhenomD. We show that the new model is faithful, with less than 1% mismatch, against a set of hybrid waveforms, except for one outlier that barely breaches this limit. The recalibrated model shows mismatches of up to $\sim 16\%$ with IMRPhenomD for high mass ratios and spins. Amplitude corrections become less significant for higher mass ratios, whereas the phase corrections leave more impact – suggesting that the former is practically irrelevant for gravitational wave data analysis in Advanced LIGO (aLIGO), Virgo and KAGRA. Comparing with a set of 219 numerical relativity waveforms, we find that the median of mismatches decreases by $\sim 4\%$ in aLIGO zero-detuned high power noise curve, and by $\sim 2\%$ with a flat noise curve. This implies a modest but notable improvement in waveform accuracy.
@@ I Introduction
Efforts to construct accurate IMR models have been based on the EOB formalism with calibrations to NR data (SEOBNR [13, 15, 16] and TEOBResumS [17, 18, 19] models), phenomenological waveform models built by combining PN and NR waveforms [20, 21, 22, 23, 24, 25], and more recently, surrogate models [26, 27, 28]. The computational cost of EOB models, however, is burdened by the need to tackle the orbital dynamics through solving a complex system of ordinary differential equations. Phenomenological models, in comparison, are equipped with closed-form expressions for the phase and amplitude of the waveforms for a given set of binary parameters, and are much faster to evaluate. The latter models contain PN-inspired phase and amplitude behaviors augmented with higher-order terms calibrated against a set of “hybrid waveforms”. The hybrids are constructed by stitching either PN (e.g., in IMRPhenomC [21]) or EOB (e.g. IMRPhenomD [25, 24], IMRPhenomXAS [23] for aligned-spin cases) waveforms with NR ones at a suitable frequency, to allow smooth transitions of phase and amplitude from the inspiral to the merger-ringdown regime. Utmost care needs to be borne when producing complete IMR waveforms by combining these parts, to limit introducing waveform systematics, which can reduce the effectiveness of tests of GR. The increasing modeling errors and computational cost of gravitational waveforms with increasing mass ratio and spins [29] have inspired the advent of reduced-order models (ROMs) and surrogate models. By fitting interpolated decomposed waveform data pieces over the binary parameter space, surrogate models can significantly accelerate NR (e.g. [30, 28, 31]) or EOB waveforms (e.g. [32, 33, 34]), while maintaining high accuracy within their parameter space of validity.
@@
We first describe some preliminaries and notations of gravitational waveforms in Sec. II. Then in Sec. III.1, we build an inspiral model containing explicit corrections due to TH at 2.5PN, 3.5PN, and 4PN orders in their phase and amplitude. The point-particle baseline for the inspiral waveform is based on an aligned-spin EOB approximant. In Sec. III.2, we describe the publicly available NR waveforms of the SXS catalog [54] used for merger and ringdown phases. In Sec. IV, we describe the construction of new hybrid waveforms by stitching together the TH-corrected inspiral and merger-ringdown parts. Following IMRPhenomD (henceforth referred to as PhenomD), we model the hybrids in the frequency domain in Sec. V and calibrate the inspiral part with augmented PN expressions, but against the new set of hybrid waveforms. Section VI discusses the faithfulness of the new model, and in Sec. VII, we compare it with PhenomD and NR data within the parameter range of calibration. We conclude in Sec. VIII with discussions and future directions of this work.
@@ III.1 Inspiral waveforms
III.1.1 Effective-one-body description of a compact binary coalescence
In the effective-one-body approach to the two-body problem in GR, originally developed by Buonanno and Damour [12], the dynamics of two compact objects of masses $m_{1}$ and $m_{2}$ and spins ${\bf\it\chi_{1}}$ and ${\bf\it\chi_{2}}$ is mapped onto the dynamics of an effective particle of mass $\mu=m_{1}m_{2}/(m_{1}+m_{2})$ and spin ${\bf\it\chi_{\ast}}$ moving in the background of a deformed Kerr geometry with mass $M=m_{1}+m_{2}$ and spin ${\bf\it\chi}_{\rm Kerr}$. The spin mapping $\{{\bf\it\chi_{1}},{\bf\it\chi_{2}}\}\to{\bf\it\chi_{\ast}}$ and the deformation of the Kerr background, parametrized by the symmetric mass ratio $\eta$, implies that the test-particle dynamics reproduces the PN-expanded dynamics of the original two-body system. Free parameters are introduced into the models that represent unknown, higher-order PN terms, or additional physical effects like corrections due to eccentricity. Such free parameters are calibrated to NR simulations. With the EOB system specified, its conservative dynamics can be described by an EOB Hamiltonian [16, 15], while the non-conservative dynamics is contained in a parametrized radiation-reaction term that is inserted in the equations of motion. This term sums over the outgoing GW modes and is calibrated to reproduce NR simulations. The combination of these two pieces describes the binary inspiral through to the merger, at which point a ringdown waveform is attached to the inspiral-merger waveform. The ringdown waveform is constructed as a linear superposition of the dominant quasinormal modes (QNMs) of the Kerr BH formed at merger [55, 53], with the amplitude and phase of each QNM mode determined by the process of stitching the ringdown part with the inspiral-merger parts.
For the purpose of creating hybrid waveforms in our work, we choose SEOBNRv2 [16], an aligned-spin EOB approximant, as the point-particle baseline. More recent versions of the SEOBNR family of waveforms are available [13] at present, but we create our model based on the former one for a direct comparison with PhenomD. We also mention here that for PhenomD, NR calibrations of SEOBNRv2 were removed before creating hybrid waveforms, referred to as SEOBv2. In this work, however, we use the original calibrated version of the EOB model since in the hybrids, stitched at a frequency much lower than the innermost stable circular orbit (ISCO), we expect little to no NR information from the inspiral approximant to be present.
@@ III.1.2 Corrections for tidal heating of black holes
with $\Psi_{\rm PP}(f)$ being the GW phase under point particle (PP) approximation that corresponds to $\mathcal{F}^{\infty}$, and $\Psi_{\rm TH}(f)$ being the phase correction due to TH. The absorbed flux can be expressed as the sum of the rates of mass increment [41] for the two BHs:
$\mathcal{F}^{\rm H}(v)=$ ${}\frac{32}{5}\eta^{2}\frac{v^{15}}{4}\sum_{i=1}^{2}\left(\frac{m_{i}}{M}\right)^{3}\left(1+3\chi^{2}_{i}\right)$ (10)
×$\times\left\{-(\hat{L}\cdot\hat{S}_{i})\chi_{i}+2\left[1+\sqrt{1-\chi^{2}_{i}}\right]\frac{m_{i}}{M}v^{3}\right\}\ ,$
where the indices $i=1,2$ denote the two BHs, and $\hat{S}_{i}$ and $\hat{L}$ are the unit vectors along the directions of the $i$th object’s spin and the orbital angular momentum, respectively.
Using this expression, one can calculate the phase contribution of tidal heating in gravitational waveforms of BBHs to be [45, 46]
$\Psi_{\rm TH}=$ ${}\frac{3}{128\eta}\left[-\frac{10}{3}H^{(\rm BBH)}_{\rm eff5}\,\log\left(v\right)\right.$ (11)
$-\frac{5}{168}v^{2}H^{(\rm BBH)}_{\rm eff5}\left(952\eta+995\right)$
$\left.+\frac{5}{3}v^{3}\log\left(v\right)(-4H^{(\rm BBH)}_{\rm eff8}+H^{(\rm BBH)}_{\rm eff5}\Psi_{\text{SO}})\right],$
$\Psi_{\text{SO}}$ is a “spin-orbit” term, given by
$\begin{split}\Psi_{\text{SO}}=&\leavevmode\nobreak\ \frac{73}{3(1+q)^{2}}\left\{q^{2}\big{(}\hat{L}\cdot\hat{S}_{1}\big{)}\chi_{1}+\big{(}\hat{L}\cdot\hat{S}_{2}\big{)}\chi_{2}\right\}\\ &+\frac{15q}{(1+q)^{2}}\left\{\big{(}\hat{L}\cdot\hat{S}_{1}\big{)}\chi_{1}+\big{(}\hat{L}\cdot\hat{S}_{2}\big{)}\chi_{2}\right\}\,.\end{split}$ (13)
We find the PN expansion of the total amplitude by expanding $\left[-E^{\prime}(v)/\mathcal{F}(v)\right]^{1/2}$ in Eq. (8) in powers of $v$ up to $v^{7}$, and separate the contribution due to TH:
@@
where $\tilde{A}_{\rm SO}$ is the spin-orbit term
@@ III.2 Numerical relativity waveforms for merger-ringdown
Simulation label
$q$
$\chi_{1}$
$\chi_{2}$
$\chi_{\rm PN}$
$N_{\rm orb}$
$F_{\rm stitch}$
1
SXS:BBH:0156
1
--0.95
--0.95
--0.79
13
0.0057
2
SXS:BBH:0151
1
--0.6
--0.6
--0.5
15
0.0054
3
SXS:BBH:0001
1
0
0
0
28
0.0041
4
SXS:BBH:0152
1
0.6
0.6
0.5
23
0.0052
5
SXS:BBH:0153
1
0.85
0.85
0.7
25
0.0052
6
SXS:BBH:0234
2
--0.85
--0.85
--0.72
28
0.0037
7
SXS:BBH:0238
2
--0.5
--0.5
--0.43
32
0.0037
8
SXS:BBH:0169
2
0
0
0
16
0.0062
9
SXS:BBH:0253
2
0.5
0.5
0.43
29
0.0046
10
SXS:BBH:2131
2
0.85
0.85
0.72
25
0.0056
11
SXS:BBH:1936
4
--0.8
--0.8
--0.71
17
0.0057
12
SXS:BBH:1418
4
--0.4
--0.5
--0.37
67
0.005
13
SXS:BBH:0167
4
0
0
0
16
0.0071
14
SXS:BBH:1417
4
0.4
0.5
0.37
80
0.0057
15
SXS:BBH:1907
4
0
0.8
0.12
21
0.0062
16
SXS:BBH:1423
8
--0.6
--0.75
--0.57
18
0.0066
17
SXS:BBH:0064
8
--0.5
0
--0.43
19
0.0066
18
SXS:BBH:0063
8
0
0
0
26
0.0065
19
SXS:BBH:0065
8
0.5
0
0.43
34
0.0063
20
SXS:BBH:1426
8
0.48
0.75
0.47
26
0.0078
@@
Figure 1: Hybrid waveforms for four different configurations in the parameter space of $q,\chi_{1},\chi_{2}.$ Inspiral waveforms are generated by adding the phase and amplitude corrections due to TH to the SEOBNRv2 model, shown in blue. NR waveforms are shown in orange, and the hybrid waveforms are shown in black dashed-dotted lines. The $x$ and $y$ axes denote time (in units of total mass) and the real part of time-domain strain, $h_{+}(t)$ (in units of $M/D_{L}$), respectively. The blue shaded areas denote the stitching regions.

### v3

@@ Abstract
Subjected to the tidal field of its companion, each component of a coalescing binary suffers a slow change in its mass (tidal heating) and spin (tidal torquing) during the inspiral and merger. For black holes, these changes are associated with their absorption of energy and angular momentum fluxes. This effect modifies the inspiral rate of the binary, and consequently, the phase and amplitude of its gravitational waveform. Numerical relativity (NR) waveforms contain these effects inherently, whereas analytical approximants for the early inspiral phase have to include them manually in the energy balance equation. In this work, we construct a frequency-domain gravitational waveform model that incorporates the effects of tidal heating of black holes. This is achieved by recalibrating the inspiral phase of the waveform model IMRPhenomD to incorporate the phase corrections for tidal heating. We also include corrections to the amplitude, but add them directly to the inspiral amplitude model of IMRPhenomD. First we demonstrate that the inclusion of the corrections, especially in the phase, confers an overall improvement in the phase agreement between the analytical inspiral model (uncalibrated SEOBNRv2) and NR data. The model presented here is faithful, with less than 1% mismatches against a set of hybrid waveforms (except for one outlier that barely breaches this limit). The recalibrated model shows mismatches of up to $\sim 16\%$ with IMRPhenomD for high mass ratios and spins. Amplitude corrections become less significant for higher mass ratios, whereas the phase corrections leave more impact – suggesting that the former is practically irrelevant for gravitational wave data analysis in Advanced LIGO (aLIGO), Virgo and KAGRA. Comparing with a set of 219 numerical relativity waveforms, we find that the median of mismatches decreases by $\sim 4\%$ in aLIGO zero-detuned high power noise curve, and by $\sim 1.5\%$ with a flat noise curve. This implies a modest but notable improvement in waveform accuracy.
@@ I Introduction
Efforts to construct accurate IMR models have been based on the EOB formalism with calibrations to NR data (SEOBNR [13, 15, 16] and TEOBResumS [17, 18, 19] families of models), phenomenological waveform models built by combining PN and NR waveforms [20, 21, 22, 23, 24, 25], and more recently, surrogate models [26, 27, 28]. EOB models generate gravitational waveforms by solving a set of ordinary differential equations. The post-adiabatic (PA) approximation has been used with notable success in the TEOBResumS [14, 29] and some of the SEOBNR [30, 31] models to solve the dynamical equations analytically, and substantially speeding up the waveform generation. Phenomenological models, in comparison, are equipped with closed-form expressions for the phase and amplitude of the waveforms for a given set of binary parameters. The latter models contain PN-inspired phase and amplitude behaviors augmented with higher-order terms calibrated against a set of “hybrid waveforms”. The hybrids are constructed by stitching either PN (e.g., in IMRPhenomC [21]) or EOB (e.g. IMRPhenomD [25, 24], IMRPhenomXAS [23] for aligned-spin cases) waveforms with NR ones at a suitable frequency, to allow smooth transitions of phase and amplitude from the inspiral to the merger-ringdown regime. Utmost care needs to be borne when producing complete IMR waveforms by combining these parts, to limit introducing waveform systematics, which can reduce the effectiveness of tests of GR. The increasing modeling errors and computational cost of gravitational waveforms with increasing mass ratio and spins [32] have inspired the advent of reduced-order models (ROMs) and surrogate models. By fitting interpolated decomposed waveform data pieces over the binary parameter space, surrogate models can significantly accelerate NR (e.g. [33, 28, 34]) or EOB waveforms (e.g. [35, 36, 37]), while maintaining high accuracy within their parameter space of validity.
@@
We first describe some preliminaries and notations of gravitational waveforms in Sec. II. Then in Sec. III.1, we build an inspiral model containing explicit corrections due to TH at 2.5PN, 3.5PN, and 4PN orders in their phase and amplitude. The point-particle baseline for the inspiral waveform is based on an aligned-spin EOB approximant. In Sec. III.2, we describe the publicly available NR waveforms of the SXS catalog [57] used for merger and ringdown phases. We investigate the phase agreement between NR data and the analytical inspiral waveforms in Sec. III.3. In Sec. IV we describe the construction of new hybrid waveforms by stitching together the TH-corrected inspiral and merger-ringdown parts. Following IMRPhenomD (henceforth referred to as PhenomD), we model the hybrids in the frequency domain in Sec. V and calibrate the inspiral part with augmented PN expressions, but against the new set of hybrid waveforms. Section VI discusses the faithfulness of the new model, and in Sec. VII, we compare it with PhenomD and NR data within the parameter range of calibration. We conclude in Sec. VIII with discussions and future directions of this work.
@@ III.1 Inspiral waveforms
III.1.1 Effective-one-body description for the point-particle waveforms
In the effective-one-body approach to the two-body problem in GR, originally developed by Buonanno and Damour [12], the dynamics of two compact objects of masses $m_{1}$ and $m_{2}$ and spins ${\bf\it\chi_{1}}$ and ${\bf\it\chi_{2}}$ is mapped onto the dynamics of an effective particle of mass $\mu=m_{1}m_{2}/(m_{1}+m_{2})$ and spin ${\bf\it\chi_{\ast}}$ moving in the background of a deformed Kerr geometry with mass $M=m_{1}+m_{2}$ and spin ${\bf\it\chi}_{\rm Kerr}$. The spin mapping $\{{\bf\it\chi_{1}},{\bf\it\chi_{2}}\}\to{\bf\it\chi_{\ast}}$ and the deformation of the Kerr background, parametrized by the symmetric mass ratio $\eta$, implies that the test-particle dynamics reproduces the PN-expanded dynamics of the original two-body system. The two primary EOB families of waveforms, namely the SEOBNR and TEOBResumS families, differ in the ways they describe the conservative Hamiltonian, the amount of PN and NR information they utilize, and the radiative parts of the waveforms [58].
For the purpose of creating hybrid waveforms in our work, we choose SEOBNRv2 [16], an aligned-spin EOB approximant, as the point-particle baseline. More recent versions of the SEOBNR family of waveforms are available [13, 31] at present, but we create our model based on the former one for a direct comparison with PhenomD. We remove the NR calibrations present in the SEOBNRv2 model, following the construction of PhenomD, since fitting to NR data may leave residues of tidal heating imprints at higher frequencies. To remove the NR calibrations, we perform the changes in the SEOBNRv2 model in LALSuite [59] as mentioned in Sec. II C in Ref. [24]. The uncalibrated version of SEOBNRv2, henceforth referred to as SEOBv2, serves as the point-particle baseline for creating the hybrid waveforms. To incorporate the corrections due to TH in BBH systems, we transform this time-domain waveform to the frequency domain to add the phase and amplitude corrections, as described in the next section.
@@ III.1.2 Corrections for tidal heating of black holes
with $\Psi_{\rm PP}(f)$ being the GW phase under point particle (PP) approximation that corresponds to $\mathcal{F}^{\infty}$, and $\Psi_{\rm TH}(f)$ being the phase correction due to TH.
The most recent expression for the dephasing $\Psi_{\rm TH}(f)$ is based on the effective worldline theory in Ref. [46], and it matches with the results in the extreme mass ratio limit [62]. In this work, we use the aforementioned results to calibrate our BBH waveform model with tidal heating. Up to 4PN order, $\Psi_{\rm TH}(f)$, or equivalently $\Psi_{\rm TH}(v)$, can be expressed as (for the (2,2) mode)
$\Psi_{\rm TH}(v)=\frac{3}{128\eta v^{5}}\sum_{n=5}^{8}\left[\delta\psi_{(n)}v^{n}+\delta\psi_{(n,l)}v^{n}\log v\right]\,.$ (10)
Explicit expressions for $\delta\psi_{(n)}$ and $\delta\psi_{(n,l)}$ are given in Sec. V of Ref [46].
Amplitude corrections for TH in the frequency domain have been derived for the first time in the current work. We find the PN expansion of the total amplitude by expanding $\left[-E^{\prime}(v)/\mathcal{F}(v)\right]^{1/2}$ in Eq. (8) in powers of $v$ up to $v^{7}$, and separate the contribution due to TH:
@@
$\tilde{A}_{\rm SO}$ is the spin-orbit term
@@
SEOBv2 is a time-domain waveform, and the phase and amplitude corrections for TH are calculated in the frequency domain. So, we have to first convert the time-domain data to the Fourier domain to apply the corrections. Since the time-domain data is finite, we use a tapered window function to minimize Gibbs artifacts, called the Planck-taper window [63]:
The frequency-domain waveform, after correcting for TH, is then converted back to the time domain. For ease of reference within this paper, we call the resulting waveform SEOBv2_Horizon. In Appendix B we list the waveform models used in this work for various purposes.
@@ III.2 Numerical relativity waveforms for merger-ringdown
III.3 Phase comparison of inspiral and NR waveforms
Figure 1: Accumulated phase difference between the inspiral and NR waveforms in 10 GW cycles of the latter prior to $M\omega_{\rm GW}=0.075$: $\phi_{\rm NR}(0)-\phi_{\rm NR}(t_{0})=20\pi$. The NR data correspond to the 20 NR waveforms used for hybrid construction, described in Table 1. (a) Top panel: variation with the mass ratio $q$, shown for two cases of the inspiral waveforms (SEOBv2), one with the tidal heating corrections and one without. The red dashed line connects the nonspinning binaries. (b) Bottom panel: same points (with the tidal heating corrections included) as in the top panel, but plotted against the effective spin parameter $\chi_{\rm eff}=(q\chi_{1}+\chi_{2})/(1+q)$. Equal-mass binaries are highlighted.
Figure 2: Difference between the accumulated phase errors (Figure 1) with and without the tidal heating corrections. (a) Top panel: the 20 NR waveforms used in Fig. 1. The subscript H (PP) refers to the SEOBv2 model with (without) the TH corrections. Spin values are reported for the binaries that produce higher phase errors compared to NR after adding the TH corrections ($|\Delta\phi|_{\rm H}-|\Delta\phi|_{\rm PP}\geq 0$), plotted in red color. Binaries with lesser phase disagreement after the corrections are plotted in green. (b) Middle panel: same as in the top panel, with 218 NR waveforms from the SXS catalog, plotted against the mass ratio. (c) Bottom panel: same as in the middle panel, plotted against the effective spin parameter.
# Simulation label $q$ $\chi_{1}$ $\chi_{2}$ $\chi_{\rm PN}$ $N_{\rm orb}$ $F_{\rm stitch}$
1 SXS:BBH:0156 1 --0.95 --0.95 --0.79 13 0.0057
2 SXS:BBH:0151 1 --0.6 --0.6 --0.5 15 0.0054
3 SXS:BBH:0001 1 0 0 0 28 0.0041
4 SXS:BBH:0152 1 0.6 0.6 0.5 23 0.0052
5 SXS:BBH:0153 1 0.85 0.85 0.7 25 0.0052
6 SXS:BBH:0234 2 --0.85 --0.85 --0.72 28 0.0037
7 SXS:BBH:0238 2 --0.5 --0.5 --0.43 32 0.0037
8 SXS:BBH:0169 2 0 0 0 16 0.0062
9 SXS:BBH:0253 2 0.5 0.5 0.43 29 0.0046
10 SXS:BBH:2131 2 0.85 0.85 0.72 25 0.0056
11 SXS:BBH:1936 4 --0.8 --0.8 --0.71 17 0.0057
12 SXS:BBH:1418 4 --0.4 --0.5 --0.37 67 0.005
13 SXS:BBH:0167 4 0 0 0 16 0.0071
14 SXS:BBH:1417 4 0.4 0.5 0.37 80 0.0057
15 SXS:BBH:1907 4 0 0.8 0.12 21 0.0062
16 SXS:BBH:1423 8 --0.6 --0.75 --0.57 18 0.0066
17 SXS:BBH:0064 8 --0.5 0 --0.43 19 0.0066
18 SXS:BBH:0063 8 0 0 0 26 0.0065
19 SXS:BBH:0065 8 0.5 0 0.43 34 0.0063
20 SXS:BBH:1426 8 0.48 0.75 0.47 26 0.0078
@@ III.3 Phase comparison of inspiral and NR waveforms
Figure 3: Hybrid waveforms for four different configurations in the parameter space of $q,\chi_{1},\chi_{2}.$ Inspiral waveforms are generated by adding the phase and amplitude corrections due to TH to the SEOBv2 model, shown in blue. NR waveforms are shown in orange, and the hybrid waveforms are shown in black dashed-dotted lines. The $x$ and $y$ axes denote time (in units of total mass) and the real part of time-domain strain, $h_{+}(t)$ (in units of $M/D_{L}$), respectively. The blue shaded areas denote the stitching regions.
In this section we investigate the phase evolution of the time-domain inspiral and NR waveforms. Construction of hybrid waveforms, as described in Sec. IV, demands that the two input waveforms be sufficiently close up to the frequency where they are stitched together. Following Hannam et al. [76, 77], we quantify this by aligning the two waveforms in phase at a certain frequency and calculating the phase difference accumulated over a number of GW cycles before this frequency. We choose to make the alignment  [^2] In Ref. [76] the choice made was $M\omega_{\rm GW}=0.1$. We choose a slightly lower frequency so that the SEOBv2 waveform without NR calibrations can be generated with sufficient accuracy. at GW frequency $M\omega_{\rm GW}=0.075$ or, equivalently, $Mf_{\rm GW}=F\approx 0.012$. The method is the following:
  1. 1.
    Given the inspiral and the NR waveforms, we calculate the times where the GW frequency reaches $F=0.012$. We label them as $t=0$ for both the waveforms.
  2. 2.
    We extract the time-domain phase of the waveforms and make the phases coincide at $t=0$.
  3. 3.
    We calculate the total phase difference that accumulates over 10 GW cycles in the NR waveform prior to $t=0$, i.e. when the phase (in rad) is $20\pi$ less than that at $t=0$. In other words, if $t=t_{0}$ is taken to be 10 GW cycles before $t=0$ in the NR waveform, then $\Delta\phi$ is defined as $\phi_{\rm INS}(t_{0})-\phi_{\rm NR}(t_{0})$.
In Fig. 1 we show the accumulated phase difference $\Delta\phi=\phi_{\rm INS}(t_{0})-\phi_{\rm NR}(t_{0})$ (in rad) between the inspiral waveform SEOBv2 and the 20 NR waveforms described in Table 1 that are used for the construction of hybrids. We show the results for both SEOBv2 and SEOBv2_Horizon. Figure 1 reports the variation with mass ratio, where we highlight the nonspinning cases. In Fig. 1 we show the phase difference as a function of the effective spin parameter $\chi_{\rm eff}=(q\chi_{1}+\chi_{2})/(1+q)$, and highlight the equal-mass binaries. The accumulated phase difference does not exceed 0.5 radians within these 10 cycles for any of these waveforms, implying that the inspiral model is suitable for hybrid construction. With the TH corrections introduced, we see from Fig 1 that the phase difference is lower in most cases, especially for high mass-ratio binaries. This is further demonstrated in Fig. 2, where we show the difference between the $|\Delta\phi|$ values before and after imposing the TH corrections. To investigate this with a larger set of NR data, we take 218 aligned-spin waveforms from the SXS catalog and repeat the procedure and plot them in Fig. 2 against the mass ratio $q$, and in Fig. 2 against $\chi_{\rm eff}$. Tidal heating introduces only small corrections in the phase; consequently, for symmetric masses and low-spin binaries, the phase error fluctuates. However, towards higher mass ratios (Figure 2) and high spins (Figure 2) a statistical behaviour emerges and, for the majority of the binaries, TH corrections contribute to a lesser phase disagreement between the inspiral and the NR waveforms.
The key takeaway from this section is that the SEOBv2_Horizon waveform is sufficiently robust for hybrid construction, with phase errors less than 0.5 radians over 10 GW cycles prior to $M\omega_{\rm GW}=0.075$. Furthermore, corrections due to tidal heating of BHs provide a statistical improvement in the phase error for asymmetric masses and highly spinning binaries, especially for anti-aligned spins.